OpenStudy (anonymous):

i left something out how do you find the prime factorization of 72 using exponents

OpenStudy (anonymous):

72 = 2 * 36 36 = 2 * 18 18 = 2 * 9 9 = 3 * 3 SO 72 = 2 * 2 * 2 * 3 * 3 2 * 2 * 2 = 2^3 3 * 3 = 3^2 SO 72 = 2^3 * 3^2

OpenStudy (anonymous):

72 /9 = 8. 9 = 3^2. 8= 2^3.